# Riccardo Giordana Pozzi and Diego Trancanelli

### v1

@@ 2.2 The superdisplacement multiplet
In ordinary dCFTs, one can identify the displacement operator associated with the broken symmetry generators of translations orthogonal to the line [49]. For supersymmetric theories, the displacement operator is accompanied by all other operators associated with the broken bulk generators, which are expected to arrange in a supermultiplet, the so-called superdisplacement multiplet.
By recalling the definition of the Wilson line (2.1), one can see that by considering infinitesimal variations it is possible to relate the broken symmetry generators $\mathbf{G}$ to defect operator insertions $\mathbb{G}(t)$ on the line [3, 50]. This procedure leads to the Ward identity [31]
@@ 3.1 Correlation functions in superspace
is the chiral distance between two points in superspace and ${\Delta}$ the conformal dimension of the superfields. The normalization $C_{\Phi}$ depends on the coupling constant of the theory and has an important physical meaning, being generally related to the Bremsstrahlung function [49]. [^6] For ABJM see also, e.g , [53, 54, 55, 56, 27] and chapters 10 and 11 of [42] for a review.
@@
with $t_{ij}\equiv t_{i}-t_{j}$. The powers of $t_{12}$ in these expressions are of course consistent with the conformal dimensions of the components of the superdisplacement (2.31).
@@ 3.2 Selection rules
Regarding the chiral-chiral case, one can study the OPE expansion by requiring that the right-hand side of the expansion respects the chirality condition and the charges. The first contribution one may expect to be exchanged is then the superprimary, as it satisfies the full shortening condition. All other contributions are instead built by acting a sufficient number of times with $\bar{Q}^{a}$, until the correct chirality condition is met. This results in (see (B.3) for the charges of $\bar{Q}^{a}$)
$L\bar{A}[j_{0}]_{j_{0}}^{(0)}\times L\bar{A}[j_{0}]_{j_{0}}^{(0)}\sim L\bar{A}[2j_{0}]_{2j_{0}}^{(0)}+\bar{Q}^{1}L\bar{A}[2j_{0}]^{(1)}_{2j_{0}+\frac{1}{2}}+\bar{Q}^{2}\bar{Q}^{1}L\bar{L}[2j_{0}]^{(0)}_{\Delta}.$ (3.21)
In this case we have explicitly written down the $\bar{Q}$-descendants given by the action of the supercharges on the corresponding superprimaries. The quantum numbers of the exchanged operators are therefore $[2j_{0},2j_{0},0]$, $[2j_{0}+1,2j_{0},0]$ and $[\Delta+1,2j_{0},0]$.
@@ 3.3.1 $f$-correlator
where $\Delta>0$ follows directly from the selection rule and the blocks sum the contributions of the conformal descendants with quantum numbers $[\Delta,0,0]$. In order to identify them, we rely on the shadow formalism [58], see [59] for a review. We consider the quadratic Casimir (2.14) acting on the operators inserted at $y_{1}$ and $y_{2}$. To this scope, we replace the generators in (2.14) with generators acting on the two coordinates, e.g. $D\rightarrow D_{\text{s}}=D_{1}+D_{2}$, $K\rightarrow K_{\text{s}}=K_{1}+K_{2}$ and so on. [^8] This is just the same as considering $\mathfrak{D}_{1,2}\propto(\mathcal{J}_{ab,1}+\mathcal{J}_{ab,2})(\mathcal{J}^{ab}_{1}+\mathcal{J}^{ab}_{2})$ where $\mathcal{J}^{ab}_{1}$ are the differential representations of the generators of the superconformal algebra and the single Casimir is $\mathfrak{C}\propto\mathcal{J}^{ab}\mathcal{J}_{ab}$ Overall, the differential operator one obtains is
@@
which is solved by the linear combination
@@
the conformal blocks have the simple expression
@@
The $G_{-1-\Delta}(z)$ term is associated with the so-called shadow contributions. Its appearance is related to the fact that $G_{-1-\Delta}(z)$ and $G_{\Delta}(z)$ have the same eigenvalues $\mathfrak{c}_{\Delta}=\mathfrak{c}_{-1-\Delta}$. Having dimension strictly less than the unitarity bound, they have no real physical interpretation, meaning that they do not belong to the physical spectrum of the theory. Therefore, we will neglect these terms in the following and keep the conformal block as defined only by $G_{\Delta}(z)$. However, with a little foresight, we can already expect that these contributions, precisely because they have the same eigenvalue under $\mathfrak{D}_{1,2}$, will be prime candidates for defining an internal (weighted) product, see 4.1.
@@ 3.3.2 $h$-correlator
We start again from the analysis of the s-channel, which is important in order to make contact with the s-channel of the $f$-correlator. Again, one expands around the insertions at $(12)$ and $(34)$, or $\chi\rightarrow 0$. One can, once more, exploit the chiral-antichiral OPE and expect $h(\chi)$ to have the same functional expansion as $f(z)$:
@@
with $G_{\Delta}$ being the same as in (3.30), this time a function of $\chi$ instead of $z$. This is essentially because both s-channels have the same chiral-antichiral OPE expansions and can be mapped to each other. This can be seen by the exchange of either the insertions $(3\leftrightarrow 4)$ or $(1\leftrightarrow 2)$, which, we stress, is not a crossing symmetry. The cross invariant ratios will be exchanged too, $z\leftrightarrow\chi$, and the final outcome is the same differential equation (3.28), this time in terms of $\chi$. It follows that the CPW coefficients will be the same as the ones of the $f$-correlator, so that
$b_{\Delta}=a_{\Delta}.$ (3.37)
@@
Now one can perform the CPW expansion, which in this case is given in terms of the conformal blocks [57]
@@ 4.1 Leading order
We start from the leading order, which is given by Wick contractions of the fields appearing in the supermatrices in (2.38) and (2.40) [61, 62, 63, 64, 4]. All cases considered in subsection 2.3 give identical functional forms.
For (3.16), one obtains
$f^{(0)}(z)=1+z^{2},\qquad h^{(0)}(\chi)=1+\chi^{2}\,,$ (4.2)
and, as expected, the two expressions have the same functional form. In fact, the operators that are exchanged in the s-channels of both 4-point functions of (3.16) are ${\mathbb{R}}\partial_{t}^{n}\bar{\mathbb{R}}$, besides the identity, while for the t-channel of the $h$-correlator they are ${\mathbb{R}}\partial_{t}^{n}{\mathbb{R}}$. All these have dimensions $\Delta^{s}_{n}=\Delta^{t}_{n}=2+n$ [^10] Here and in the following, $\Delta^{t}_{n}$ and $\gamma_{n}^{t}$ refer exclusively to the t-channel of the $h$-correlator, i.e. the chiral-chiral channel.
@@
while for the t-channel, only even $n$ contribute:
@@ 4.2 Next-to-leading order
Similarly, one can find an analogous expression for $h^{(1)}$ by recalling that in the s-channels the two functions have the same functional form. Hence, by replacing $z\rightarrow\chi$, one obtains the expression for $h^{(1)}(\chi)$ , up to the sign in the logarithm [31]
@@
In order to state the final answer one can proceed by absorbing one of the coefficient in the expansion parameter $\epsilon$ and, in analogy with the $SU(1,1|3)$ case [31], we also change the overall sign. [^11] In fact, the discussion in this section is true regardless of the overall sign of the ansatz. This sign change was justified in [31] by the comparison with the holographic dual. Hence, by reabsorbing $q_{-2}$ into the definitions of $\epsilon$ and expressing the ratio of the two free parameters as
@@ Acknowledgements
We are indebted to Gabriel Nagaoka for collaboration at the initial stages of this work and to Lorenzo Bianchi and Carlo Meneghelli for discussions. DT is supported in part by the INFN grant Gauge and String Theory (GAST) and would like to thank FAPESP’s partial support through the grant 2019/21281-4.

### v2

@@ 2.2 The superdisplacement multiplet
In ordinary dCFTs, one can identify the displacement operator associated with the broken symmetry generators of translations orthogonal to the line [49, 50, 51]. For supersymmetric theories, the displacement operator is accompanied by all other operators associated with the broken bulk generators, which are expected to arrange in a supermultiplet, the so-called superdisplacement multiplet. [^3] Parts of this derivation of the superdisplacement multiplet have been also obtained in [52].
By recalling the definition of the Wilson line (2.1), one can see that by considering infinitesimal variations it is possible to relate the broken symmetry generators $\mathbf{G}$ to defect operator insertions $\mathbb{G}(t)$ on the line [3, 53]. This procedure leads to the Ward identity [54, 31]
@@
This analysis is consistent with [43], where our superdisplacement multiplet is also considered as part of their general classification of superconformal defects.
@@ 3.1 Correlation functions in superspace
is the chiral distance between two points in superspace and ${\Delta}$ the conformal dimension of the superfields.
@@
with $t_{ij}\equiv t_{i}-t_{j}$. The powers of $t_{12}$ in these expressions are of course consistent with the conformal dimensions of the components of the superdisplacement (2.31). The normalization $C_{\Phi}$ depends on the coupling constant of the theory and, for the superdisplacement multiplet, it has an important physical meaning, being generally related to the Bremsstrahlung function [50]. [^8] For ABJM see also, e.g , [57, 58, 59, 60, 27] and chapters 10 and 11 of [42] for a review.
@@ 3.2 Selection rules
Regarding the chiral-chiral case, one can study the OPE expansion by requiring consistency under the chirality condition and the charges. In particular, if one considers the OPE for $[j_{0}]_{j_{0}}^{(0)}\times[j_{0}]_{j_{0}}^{(0)}$ where $[j_{0}]_{j_{0}}^{(0)}$ is the superprimary of $LA[j_{0}]_{j_{0}}^{(0)}$, the first contribution one expects to be exchanged is again a superprimary, as it satisfies the full shortening condition, with quantum numbers $[2j_{0}]_{2j_{0}}^{(0)}$. Notice, however, that it is not the full $LA[2j_{0}]_{2j_{0}}^{(0)}$ multiplet that takes part in the OPE, but just one of its component (in this case the superprimary), with all its conformal family. Analogously, all other contributions that can be exchanged are constrained by chirality. These are built by acting a sufficient number of times with $\bar{Q}^{a}$ on a given superprimary, until the correct chirality condition and charge consistency are met. The result for the supermultiplets is schematically given by
$L\bar{A}[j_{0}]_{j_{0}}^{(0)}\times L\bar{A}[j_{0}]_{j_{0}}^{(0)}\sim L\bar{A}[2j_{0}]_{2j_{0}}^{(0)}+L\bar{A}[2j_{0}]^{(1)}_{2j_{0}+\frac{1}{2}}+L\bar{L}[2j_{0}]^{(0)}_{\Delta}.$ (3.21)
Then, when considering the superprimary OPE, one obtains (see (B.3) for the charges of $\bar{Q}^{a}$)
$[j_{0}]_{j_{0}}^{(0)}\times[j_{0}]_{j_{0}}^{(0)}\sim[2j_{0}]_{2j_{0}}^{(0)}+\bar{Q}^{1}[2j_{0}]^{(1)}_{2j_{0}+\frac{1}{2}}+\bar{Q}^{2}\bar{Q}^{1}[2j_{0}]^{(0)}_{\Delta},$ (3.22)
where we have explicitly written down the $\bar{Q}$-descendants given by the action of the supercharges on the corresponding superprimaries. The quantum numbers of the exchanged conformal primaries are therefore $[2j_{0},2j_{0},0]$, $[2j_{0}+1,2j_{0},0]$ and $[\Delta+1,2j_{0},0]$.
@@ 3.3.1 $f$-correlator
where $\Delta>0$ follows directly from the selection rule and the blocks sum the contributions of the superconformal descendants with quantum numbers $[\Delta,0,0]$. In order to identify them, we rely on the shadow formalism [62], see [63] for a review. We consider the quadratic Casimir (2.14) acting on the operators inserted at $y_{1}$ and $y_{2}$. To this scope, we replace the generators in (2.14) with generators acting on the two coordinates, e.g. $D\rightarrow D_{\text{s}}=D_{1}+D_{2}$, $K\rightarrow K_{\text{s}}=K_{1}+K_{2}$ and so on. [^9] This is just the same as considering $\mathfrak{D}_{1,2}\propto(\mathcal{J}_{ab,1}+\mathcal{J}_{ab,2})(\mathcal{J}^{ab}_{1}+\mathcal{J}^{ab}_{2})$ where $\mathcal{J}^{ab}_{1}$ are the differential representations of the generators of the superconformal algebra and the single Casimir is $\mathfrak{C}\propto\mathcal{J}^{ab}\mathcal{J}_{ab}$ Overall, the differential operator one obtains is
@@
which is solved by the linear combination [^10] An initial derivation of these conformal blocks is contained in [52].
@@
the superconformal blocks have the simple expression
@@
The $G_{-1-\Delta}(z)$ term is associated with the so-called shadow contributions. Its appearance is related to the fact that $G_{-1-\Delta}(z)$ and $G_{\Delta}(z)$ have the same eigenvalues $\mathfrak{c}_{\Delta}=\mathfrak{c}_{-1-\Delta}$. Having dimension strictly less than the unitarity bound, they have no real physical interpretation, meaning that they do not belong to the physical spectrum of the theory. So, consistently with the OPE, one sets $c_{1}=0$ and $c_{2}=1$. However, with a little foresight, we can already expect that these contributions, precisely because they have the same eigenvalue under $\mathfrak{D}_{1,2}$, will be prime candidates for defining an internal (weighted) product, see 4.1.
Moreover, is worth noticing that the superblocks can be consistently decomposed in terms of 1-dimensional conformal blocks $\tilde{g}_{\Delta}(z)=(-z)^{\Delta}\,{}_{2}F_{1}(\Delta,\Delta,2\Delta;z)$ as follows
$G_{\Delta}(z)=\tilde{g}_{\Delta}(z)+\frac{\Delta}{2+2\Delta}\tilde{g}_{\Delta+1}(z)+\frac{\Delta^{2}}{4(3+4\Delta(2+\Delta))}\tilde{g}_{\Delta+2}(z)\,.$ (3.33)
@@ 3.3.2 $h$-correlator
We start again from the analysis of the s-channel, which is important in order to make contact with the s-channel of the $f$-correlator. Again, one expands around the insertions at $(12)$ and $(34)$, or $\chi\rightarrow 0$. One can, once more, exploit the chiral-antichiral OPE and expect $h(\chi)$ to have the functional expansion
@@
with $G_{\Delta}=\chi^{\Delta}\,{}_{2}F_{1}(\Delta,\Delta,2\Delta+2;\chi)$ being obtained by solving (3.31) and by imposing consistency with the OPE expansions. In this respect, the s-channel OPE coefficients of both correlators are specified by the 3-point coefficients
$a_{\Delta}=c_{\mathbb{R}\bar{\mathbb{R}}\mathcal{O}_{\Delta}}c_{\mathbb{R}\bar{\mathbb{R}}\mathcal{O}_{\Delta}},\qquad b_{\Delta}=c_{\mathbb{R}\bar{\mathbb{R}}\mathcal{O}_{\Delta}}c_{\bar{\mathbb{R}}\mathbb{R}\mathcal{O}_{\Delta}},$ (3.39)
where $\mathcal{O}_{\Delta}$ denotes the exchanged operator of conformal dimension $\Delta$. As it can be seen by applying a parity transformation [31], the two coefficients are related by
$a_{\Delta}=(-1)^{s_{\mathcal{O}_{\Delta}}}b_{\Delta},$ (3.40)
where $s_{\mathcal{O}_{\Delta}}$ stands for the charge of the exchanged operator under parity. Taking, for example, operators schematically defined as $\mathcal{O}_{2+n}\sim\mathbb{R}\partial_{t}^{n}\bar{\mathbb{R}}$, see section 4, the corresponding charge is $s_{\mathcal{O}_{2+n}}=n$. In passing, let us also notice that (3.40) depends solely on the quantum number of the exchanged operator under parity, and thus it remains valid even when perturbations are considered.
@@
Now we can perform the CPW expansion and, again, we rely on the analysis of subsection 3.2 regarding the chiral-chiral channel. In this case, we have seen that only a single component of each supermultiplets, with all its conformal family, participates the OPE. It follows that, in this case, the expansion is given in terms of the 1-dimensional conformal blocks [61]
@@ 4 Analytic bootstrap
By general arguments, the parameter $\epsilon$ scales like the inverse string tension $1/T\sim 1/\sqrt{\lambda}$ and is therefore small at strong coupling in the gauge theory, as stated above. The precise mapping depends on the specific ${\cal N}=4$ theory (and, therefore, coupling $\lambda$) one considers and can only be established after an explicit comparison with the Witten diagram computation, not being fixed by symmetry.
To emphasize the physical meaning of this small parameter, we derive below an explicit relation between $\epsilon$ and $C_{\Phi}$ (or, equivalently, between $\epsilon$ and the Bremsstrahlung function) following the ideas in [65, 66] based on the computation of the Zamolodchikov metric of the defect conformal manifold. [^12] We thank Nadav Drukker for suggesting this computation.
@@ 4.1 Leading order
We start from the leading order, which is given by Wick contractions of generalized free fields (GFF) [67, 68, 69, 70, 4]. For example, the 4-point correlation function $\langle\mathbb{R}(t_{1})\bar{\mathbb{R}}(t_{2}){\mathbb{R}}(t_{3})\bar{\mathbb{R}}(t_{4})\rangle$ factorizes in this limit as products of 2-point correlation functions
$\langle\mathbb{R}(t_{1})\bar{\mathbb{R}}(t_{2})\rangle\langle{\mathbb{R}}(t_{3})\bar{\mathbb{R}}(t_{4})\rangle+\langle\mathbb{R}(t_{1})\bar{\mathbb{R}}(t_{4})\rangle\langle{\bar{\mathbb{R}}}(t_{2})\mathbb{R}(t_{3})\rangle=C_{\Phi}^{2}\Bigl{(}\frac{1}{t_{12}^{2}t_{34}^{2}}+\frac{1}{t_{14}^{2}t_{32}^{2}}\Bigr{)}=\frac{C_{\Phi}^{2}}{t_{12}^{2}t_{34}^{2}}(1+z^{2}),$ (4.2)
where (3.9) and the explicit expression (3.13) for $z$ have been used. For (3.16), one obtains
$f^{(0)}(z)=1+z^{2},\qquad h^{(0)}(\chi)=1+\chi^{2}\,.$ (4.3)
As expected, the two expressions have the same functional form. In fact, the operators that are exchanged in the s-channels of both 4-point functions of (3.16) are ${\mathbb{R}}\partial_{t}^{n}\bar{\mathbb{R}}$, besides the identity, while for the t-channel of the $h$-correlator they are ${\mathbb{R}}\partial_{t}^{n}{\mathbb{R}}$. All these have dimensions $\Delta^{s}_{n}=\Delta^{t}_{n}=2+n$ [^13] Here and in the following, $\Delta^{t}_{n}$ and $\gamma_{n}^{t}$ refer exclusively to the t-channel of the $h$-correlator, i.e. the chiral-chiral channel.
@@
and, from the discussion of subsection 3.3.2, one finds $b^{(0)}_{n}=(-1)^{n}\,a_{n}^{(0)}$. For the t-channel, only even $n$ contribute:
@@ 4.2 Next-to-leading order
Similarly, one can find an analogous expression for $h^{(1)}$ by recalling that in the s-channels the two functions have the same functional form. Hence, by replacing $z\rightarrow\chi$, one obtains the expression for $h^{(1)}(\chi)$ , up to the sign in the logarithm [^14] The sign is changed in order to require consistency with the first-order expansion of the expression (3.38), see also [31]. By doing the expansion explicitly two contributions given by $\log(\chi)$ and $\log(1-\chi)$ factorize. Equivalently, one can consider absolute values in the arguments as done in [4].
@@
In order to state the final answer one can proceed by absorbing one of the coefficient in the expansion parameter $\epsilon$ and, in analogy with the $SU(1,1|3)$ case [31], we also change the overall sign.151515 In fact, the discussion in this section is true regardless of the overall sign of the ansatz. This sign change was justified in [31] by the comparison with the holographic dual. Hence, by reabsorbing $q_{-2}$ into the definitions of $\epsilon$ and expressing the ratio of the two free parameters as
@@ 4.5 Comments on the holographic dual
4.6 Relation between $\epsilon$ and $C_{\Phi}$
It is possible to establish a relation between the expansion parameter $\epsilon$ and the normalization of the 2-point correlation functions $C_{\Phi}$ or, equivalently, the Bremsstrahlung function of the theory. This was first proposed in [65] and then extended to higher orders at strong coupling in [66].
The idea is to consider the defect conformal manifold associated with exactly marginal deformations of our 1/2 BPS operators. This manifold is $\mathbb{CP}^{1}$ with Zamolodchikov metric given by
$g_{z\bar{z}}=\langle\mathbb{R}(0)\bar{\mathbb{R}}(1)\rangle=C_{\Phi}\delta_{z\bar{z}}\,,$ (4.54)
which is conformally related to the line element
$ds^{2}_{\mathbb{CP}^{1}}=\frac{4C_{\Phi}}{(1+|z|^{2})^{2}}dzd\bar{z}.$ (4.55)
The corresponding Riemann tensor is
$R_{z\bar{z}z\bar{z}}=-\frac{1}{2}g_{{z\bar{z}}}g_{{z\bar{z}}}\mathcal{R}_{\mathbb{CP}^{1}},\qquad\,R_{zz\bar{z}\bar{z}}=0,$ (4.56)
with Ricci scalar
$\mathcal{R}_{\mathbb{CP}^{1}}=\frac{2}{C_{\Phi}}.$ (4.57)
Now, as suggested in [65], the Riemann tensor can be also interpreted as the 4-point correlation functions of the primary operator, by extending the Zamolodchikov metric beyond the flat space approximation. There are the two possible orderings seen above, so that one has to consider
$t^{2}_{12}t^{2}_{34}\langle\mathbb{R}(t_{1})\bar{\mathbb{R}}(t_{2})\mathbb{R}(t_{3})\bar{\mathbb{R}}(t_{4})\rangle$ $=g_{z\bar{z}}g_{z\bar{z}}f(z)=C_{\Phi}^{2}f(z),$ (4.58)
$t^{2}_{12}t^{2}_{34}\langle\mathbb{R}(t_{1})\bar{\mathbb{R}}(t_{2})\bar{\mathbb{R}}(t_{3})\mathbb{R}(t_{4})\rangle$ $=g_{z\bar{z}}g_{\bar{z}z}h(\chi)=C_{\Phi}^{2}h(\chi).$ (4.59)
Following the analysis in [65] and identifying our functions with their functions as $f=2K_{1}=-2K_{2}$ and $h=2H_{1}=-2H_{2}$, one finds $R_{zz\bar{z}\bar{z}}=0$, consistently with (4.56) above, and $R_{z\bar{z}z\bar{z}}=2g_{z\bar{z}}g_{z\bar{z}}\mathcal{R}$ with
[…]
Evaluating this with $f^{(1)}$ and $h^{(1)}$, one obtains
$\mathcal{R}=2\pi^{2}\epsilon.$ (4.61)
Comparing with the Riemann tensor in (4.56), one finds the desired relation
$\epsilon=-\frac{1}{4\pi^{2}C_{\Phi}},$ (4.62)
with the minus sign being consistent with having absorbed another minus sign in $\epsilon$ in the Ansatz (see discussion in footnote 15).
If one knows the Bremsstrahlung function of the theory at strong coupling, this relation can be used to determine $\epsilon$ in a way which is independent of the comparison with the holographic computation in terms of Witten diagrams. This was done in [65] for the 1/2 BPS Wilson loops of ${\cal N}=4$ super Yang-Mills and of the ABJM theory. Unfortunately, such computation of the Bremsstrahlung function has not been performed yet for ${\cal N}=4$ Chern-Simons-matter theories, so that one cannot get the precise expression of $\epsilon$. However, the leading term of the Bremsstrahlung function scales generically like $\sqrt{\lambda}$, so that $\epsilon\sim 1/\sqrt{\lambda}$, as stated at the beginning of this section.
@@ Acknowledgements
We are indebted to Gabriel Nagaoka for collaboration at the initial stages of this work and to Lorenzo Bianchi, Nadav Drukker and Carlo Meneghelli for discussions and suggestions. DT is supported in part by the INFN grant Gauge and String Theory (GAST) and would like to thank FAPESP’s partial support through the grant 2019/21281-4.